OpenStudy (anonymous):

How to solve 5 + 7 - 3x = 4 I need an answer today, by the way

OpenStudy (shinalcantara):

combine all the constants -3x = 4 - 5 - 7 simplify. can you do that?

OpenStudy (shinalcantara):

simplify the terms at the right first

OpenStudy (igreen):

5 + 7 - 3x = 4 Add: 12 - 3x = 4 Subtract 12 to both sides: -3x = -8 Can you divide -8 / -3? @GreenNerdGirl

OpenStudy (igreen):

Yeah, that's your answer..do you have choices though?

OpenStudy (igreen):

It could just be a fraction as the answer..

OpenStudy (anonymous):

Well at first I thought it was 8/3, but I knew something was wrong with it, but I was just checking to make sure

OpenStudy (igreen):

8/3 is correct.. -8/-3 is the same thing as 8/3.

OpenStudy (anonymous):

Thanks for the help @iGreen and @shinalcantara